8. For which values of does the boundary-value problem y – 2y+(1+1)y=0; y(0)=0, y(1)=0 have a nontrivial solution?
Answer #1

y!- 2y+11+1) 4 = 0, 4(0) = 0 = 419 The A-E is m²- 2m + (1+d] =0 = m = 2:19-41142) - 2 Case I: s=o Then y= (1 + (2x) ex 410)